# SoK: Payment Channel Networks

### v1

@@ I Introduction
TABLE I: Comparison of Pathfinding and Routing Protocols in PCNs. A $\checkmark$ implies that the protocol satisfies the property and an implies that the routing protocol does not satsify the property
@@
Auroch [35] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
@@ III Virtual channels
Jia et al. [44] UTXO and TC Limited $\checkmark$
@@
Motivation: We have already described how PCNs offer a solution to the blockchain scalability problem by facilitating instantaneous payments in an off-chain manner. PCNs however, have few limitations: 1) The path generated in a PCN to process a payment between a sender and a receiver usually contains one or several intermediate nodes. All the nodes along the path are required to be available to route the payment until the receiver. This may not be possible at all times since, nodes can be disconnected from the network due to various reasons such as power outages, DoS attacks, etc. In a virtual channel, once the coins of the intermediary(ies) are locked, it can choose to remain offline without affecting the transaction processing. 2) Each node along the path from the sender to the receiver charges a fees called as routing fees/forwarding fees for forwarding the payment to the next node along the path. The sender has to pay all the nodes their respective forwarding fee along the path. This increases the overall cost of routing the payment. Virtual channels provide a solution to this problem, since they do not charge fees on a per transaction basis. Instead, the fees is charged based on a time epoch basis (the time for which the coins of the intermediary/intermediaries are locked in the virtual channel) or a fixed fee can be agreed in advance by all the parties involved in the virtual channel construction. 3) The time taken to route a payment from a sender to a receiver is directly proportional to the path length. A longer path would result in an increase in the overall pahtfinding and routing time. Virtual channels do not process payments through a path comprising of several intermediate nodes, hence the overall transactions’ pathfinding and routing time does not depend on the number of intermediate nodes between the sender and the receiver.
Virtual channels should not add an unnecessary burden to users, and should mirror the operations of payment channels as closely as possible, with the added benefit of no on-chain transactions at all, while maintaining comparable security/privacy properties. We now give the properties desired from an ideal virtual channel, and compare the works in this area on the extent to which they achieve these properties.
@@
6) Recursive: A virtual channel is said to be recursive or hierarchical if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by recursively constructing virtual channels on top of existing virtual channels from sender to receiver. We give an example of a recursive virtual channel in Figure 3(c).
7) Multihop: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by recursively constructing a single virtual channel from the sender to the receiver, without establishing virtual channel between any pair of intermediate nodes along the path. We give an example of a recursive virtual channel in Figure 3(b).
@@ IV State channels
Motivation: State channels allow the already existing payment channels to facilitate execution of arbitrary applications such as games, e-commerce, etc. This leads to the PCN becoming more versatile. State channels should not add an unnecessary burden to users in terms of programming/scripting requirements, with the added benefit of facilitating the execution of any decentralized application between them, while maintaining comparable security properties and should ideally be able to resolve disputes without accessing the blockchain. We now give the properties desired from an ideal state channel, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of various state channel protocols in Table III. We describe the metrics used for comparison below.
@@ V Tumblers
A payment channel hub should should be able to facilitate anonymous payments between a pair of sender/receiver with the added benefit of ensuring that the payment is sent to the receiver or it does not go through at all, while maintaining privacy properties. We now give the properties desired from an ideal payment channel hub, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of off-chain payment channel hubs in the Table IV. We describe the comparison metrics below:
@@
4) Time epochs: The tumbler processes transactions in discrete fragments of time called epochs, i.e., transactions only take place during a time epoch. A payment channel hub that supports transaction processing based on time epochs is not efficient. Since nodes would ideally need to wait until a new time epoch begins to process payments again. The duration between time epochs depends on several factors such as the unit of the time epoch (absolute time or relative time measured in terms of block height), number of users that will most likely involved in the next time epoch, etc.
5) Dynamic corruption: The payment channel hub should be able to handle the deviation of any party from the protocol at any point during its execution. i,e the set of adversaries should not be static during the protocol execution.
@@ VI Rebalancing
Motivation:Rebalancing in payment channel networks involves redistributing the balance within an existing payment channel or across multiple channels to maintain or improve the network’s liquidity and usability. This is done to ensure that both parties in the channel have sufficient funds on their respective sides to send and receive payments. Rebalancing is crucial to payment channel networks. If the link weights of nodes become zero/depleted as a result of being involved in transactions, it will prevent them from taking part in further payments until those link weights are replenished. Such nodes are called as dormant nodes in the network. A PCN which has a large number of such dormant nodes experiences an overall reduction in transaction throughput and an overall reduction in liquidity. Hence, the link weights of the nodes in the PCN need to be rebalanced to prevent failure of transactions due to lack of liquidity. A rebalancing protocol should be be able to replenish the link weight of dormant nodes in the PCN irrespective of the PCNs topology, while not having to employ any trusted entity and should not compromise privacy of nodes. We now give the properties desired from an ideal virtual channel, and compare the works in this area on the extent to which they achieve these properties.
We give a qualitative comparison of rebalancing protocols in Table V. We describe the metrics as follows:
@@ VII-A Gaps in pathfinding and routing for PCNs
1) Super node liquidity validation in PCNs: In payment channel networks, pathfinding and routing, which is finding a path between the sender and the receiver that can route the amount specified by the sender has attracted significant attention from the research community. Several elegant and protocols have been proposed to solve this problem [28, 89, 21, 24, 25] . In a PCN, nodes in the network store information about other nodes and their payment channels in their local storage called gossip table/gossip network. The amount of information a node has in its gossip table is a function of the number of payment channels it has open in the network. It is also a function of how much information nodes in the PCN are willing to share with their neighbors (nodes with which they have open payment channels) as a part of the gossiping protocol. The amount of information a node can attain using gossiping may be limited due to the following reasons: a node choose to share information about the network partially, a node can also choose to be anonymous, or a node’s own information can sometimes be limited due to the lack of having open payment channels.
@@
One of the main problems with the super nodes is that the sender has no way to know whether the super node possess the local balance (liquidity) to route the payment of the sender. The local balance of a super node or any node in a PCN is a private value and should not be known to any node except for its immediate neighbor. Currently, if a super node does not have enough liquidity to route the payment of a sender, the transaction fails and it has to be retried by the sender. In $\mathsf{LN}$, one of the most widely used PCN, this is a significant problem. Sometimes the sender might have to keep retrying for $\approx$ 1 hour to have a successful payment [90]. The main goal of PCNs is to facilitate instantaneous payments and these transaction retries render such payments almost impossible. It might greatly benefit the sender if the it has a mechanism to validate whether a super node has enough liquidity (balance) to route its payment without violating any privacy concerns in the network.
@@
2) Channel verification in a PCN: To be a part of any PCN, nodes will open payment channels with other nodes in the network. Two nodes open a payment channel between them by posting a transaction to the blockchain. This transaction can be posted on the blockchain (as in UTXO based blockchains like BTC) or this transaction can be posted as a function call to an existing smart contract (as in blockchains that support smart contracts like ETH). In the most popular PCN, $\mathsf{LN}$, the channel verification procedure is very inefficient. It involves the node who intends to verify a channel to request the block in which the channel opening transaction has been included, verifying whether the transaction has been successfully executed by the validator/miner and finally verifying if the channel opening transaction corresponds to a 2-2 multi signature address on the blockchain. The verifier performing these steps is inefficient since all these steps will have already been performed by the miner. According to the current verification process any node would need to perform any these steps to all the payment channels its neighbors have open. Consider the following scenario: a node $j$ in the PCN has payment channels open with $m$ nodes in the PCN. Furthermore, assume that each neighbor opens $k$ new channels for each time epoch. For this example, assume that the time epoch has been set to 1 block. This would mean that the total number of channels the node $j$ has to verify a total of $m*k$ channels. This requires the node $j$ to access the blockchain a total of $m*k$ times to perform channel verification. Finding a way to do this without blockchain access and in a blockchain agnostic manner in a hard research challenge.
@@ VII-B Gaps in Tumblers
1) Privacy against aborts: In a given a transaction, the tumbler should not be able to correctly identify the sender/receiver pair involved. Currently, there is only one construction [59] that guarantees this. However, this construction works only for blockchains that support Turing complete programming languages. Currently, there is no such construction for UTXO based blockchains that guarantees privacy against aborts. This is hard because if the sender aborts a transaction, the tumbler will know which receiver did not receive the payment. This is hard because, a tumbler mixes multiple transactions together to ensure unlikability among payments. Also, a tumbler should enforce atomicity so that the payment either goes through completely or does not go through at all. In the case of an abort, an adversary who controls the tumbler might be able to deduce which receiver amongst the ones involved did not receive a payment and will be able to link the corresponding transaction amount to that receiver.
@@
2) Time epochs: A tumbler should be able to facilitate payments without having a predefined time interval or epoch. Currently, the majority of tumbler constructions process payments in discrete epochs.To increase their anonymity sets. An anonymity set refers to a set of users (senders/receivers) whose identity is indistinguishable from each other. The larger is the size of this set, more transactions will be mixed together, increasing privacy of sender/receiver in a transaction. When the interval between two consecutive time epochs is sufficiently large, it will promote more users to take part in the transactions within a time epoch, which improves privacy. Not having time epochs, would mean the tumbler can process transactions at any instant, resulting in smaller anonymity sets and hence could possibly weaken privacy. The trade off is between privacy and flexibility. Processing transactions in discrete time epochs results in increased privacy and reduced flexibility. However, the tumbler’s ability to process transactions at any time instant might result in smaller anonymity sets, reducing privacy.
@@
3) Dynamic corruption: A tumbler should ideally be resistant to dynamic corruption. This means the protocol should be resilient to parties deviating from their ideal execution during the beginning, progress, and ending of the protocol’s execution. Current tumbler constructions provide support for static corruption, which means that an adversary decides beforehand the parties they intend to corrupt and that cannot change during the protocol execution, which may be unrealistic from a practical standpoint. Security against dynamic corruption is hard to achieve because the tumbler mixes several transactions together to preserve anonymity. Since the tumbler should not be able to link transaction amount to a sender/receiver pair (ideally), if a party gets corrupted during the protocol’s execution, the tumbler cannot correctly map the corrupted party to a particular transaction. If the sender/receiver becomes malicious during the protocol’s execution, ideally, that transaction must be aborted to prevent financial loss to the honest party. However, the tumbler should not be able to link a transaction to a sender/receiver pair, making security against dynamic corruption a hard research problem.

### v2

@@ I Introduction
TABLE I: Comparison of Pathfinding and Routing Protocols in PCNs.
@@
Auroch [35] $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ (dynamic)
@@ III Virtual channels
Jia et al. [44] UTXO and TC Limited $\checkmark$ (fixed)
@@
Motivation: We have already described how PCNs offer a solution to the blockchain scalability problem by facilitating instantaneous payments in an off-chain manner. PCNs however, have few limitations: 1) The path generated in a PCN to process a payment between a sender and a receiver usually contains one or several intermediate nodes. All the nodes along the path are required to be available to route the payment until the receiver. This may not be possible at all times since, nodes can be disconnected from the network due to various reasons such as power outages, DoS attacks, etc. In a virtual channel, once the coins of the intermediary(ies) are locked, it can choose to remain offline without affecting the transaction processing. 2) Each node along the path from the sender to the receiver charges a fees called as routing fees/forwarding fees for forwarding the payment to the next node along the path. The sender has to pay all the nodes their respective forwarding fee along the path. This increases the overall cost of routing the payment. Virtual channels provide a solution to this problem, since they do not charge fees on a per transaction basis. Instead, the fees is charged based on a time epoch basis (the time for which the coins of the intermediary/intermediaries are locked in the virtual channel) or a fixed fee can be agreed in advance by all the parties involved in the virtual channel construction. 3) The time taken to route a payment from a sender to a receiver is directly proportional to the path length. A longer path would result in an increase in the overall pathfinding and routing time. Virtual channels do not process payments through a path comprising of several intermediate nodes, hence the overall transactions’ pathfinding and routing time does not depend on the number of intermediate nodes between the sender and the receiver.
Virtual channels should not add an unnecessary burden to users, and should mirror the operations of payment channels as closely as possible, with the added benefit of no on-chain transactions at all, while maintaining comparable security/privacy properties. We now give the properties desired from an ideal virtual channel, and compare the works in this area in Table II on the extent to which they achieve these properties.
@@
6) Recursive: A virtual channel is said to be recursive or hierarchical if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by recursively constructing virtual channels on top of existing virtual channels from sender to receiver. We give an example of a recursive virtual channel in Figure 3(c). 7) Multihop: A virtual channel is said to be multihop if it can facilitate payments between a sender and a receiver across a path comprising of several intermediate nodes by constructing a single virtual channel from the sender to the receiver, without establishing virtual channel between any pair of intermediate nodes along the path. We give an example of a multihop virtual channel in Figure 3(b).
@@ IV State channels
Motivation: State channels allow the already existing payment channels to facilitate execution of arbitrary applications such as games, e-commerce, etc. This leads to the PCN becoming more versatile. State channels should be deployable on any blockchain regardless of the underlying programming/scripting requirements and should be able to facilitate the execution of any decentralized application between the parties involved, while maintaining comparable security properties (in comparison to payment channels) and should ideally be able to resolve disputes without accessing the blockchain. We now give the properties desired from an ideal state channel, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of various state channel protocols in Table III. We describe the metrics used for comparison below.
@@ V Tumblers
A payment channel hub should be able to facilitate payments between a sender and receiver, who do not have a payment channel open between them such that the hub cannot link a given transaction amount to a particular sender/receiver pair. Furthermore, the payment channel hub should also guarantee the fundamental property of atomicity (ensuring that the payment is sent to the receiver or it does not go through at all). We now give the properties desired from an ideal payment channel hub, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of off-chain payment channel hubs in the Table IV. We describe the comparison metrics below:
@@
4) Time epochs: The tumbler processes transactions in discrete fragments of time called epochs, i.e., transactions only take place during a time epoch. A payment channel hub that supports transaction processing based on time epochs is not efficient because nodes cannot process payments between time epochs, which can be of arbitrary length. The duration between time epochs depends on several factors such as the unit of the time epoch (absolute time or relative time measured in terms of block height), number of users that will most likely involved in the next time epoch, etc. 5) Dynamic corruption: The payment channel hub should be able to handle the deviation of any party from the protocol at any point during its execution. i,e the set of adversaries should not be static during the protocol execution.
@@ VI Rebalancing
Motivation:Rebalancing in payment channel networks involves redistributing the balance within an existing payment channel or across multiple channels to maintain or improve the network’s liquidity and usability. This is done to ensure that both parties in the channel have sufficient funds on their respective sides to send and receive payments. Rebalancing is crucial to payment channel networks. If the link weights of nodes become zero/depleted as a result of being involved in transactions, it will prevent them from taking part in further payments until those link weights are replenished. Such nodes are called as dormant nodes in the network. A PCN which has a large number of such dormant nodes experiences an overall reduction in transaction throughput and an overall reduction in liquidity. Hence, the link weights of the nodes in the PCN need to be rebalanced to prevent failure of transactions due to lack of liquidity. A rebalancing protocol should be be able to replenish the link weight of dormant nodes in the PCN irrespective of the PCNs topology, while not having to employ any trusted entity and should not compromise privacy of nodes. We now give the properties desired from an ideal virtual channel, and compare the works in this area on the extent to which they achieve these properties. We give a qualitative comparison of rebalancing protocols in Table V. We describe the metrics as follows:
@@ VII-A Gaps in pathfinding and routing for PCNs
1) Super node liquidity validation in PCNs: A super node is a highly connected node (it has payment channels open with numerous other nodes in the network) and also has a high liquidity in those channels. These super nodes aid other nodes in pathfinding and routing of payments. These super nodes are called as a trampoline nodes in $\mathsf{LN}$ [89, 90, 91, 92]. Several other works in the literature refer to super nodes as routing node, routing helper, landmark node, routers, etc [25, 24, 27, 37, 33]. One of the main problems with the super nodes is that the sender has no way to know whether the super node possess the local balance (liquidity) to route the payment of the sender. The local balance of a super node or any node in a PCN is a private value and should not be known to any node except for its immediate neighbor. Currently, if a super node does not have enough liquidity to route the payment of a sender, the transaction fails and it has to be retried by the sender. In $\mathsf{LN}$, one of the most widely used PCN, this is a significant problem. Sometimes the sender might have to keep retrying for $\approx$ 1 hour to have a successful payment [93]. The main goal of PCNs is to facilitate instantaneous payments and these transaction retries render such payments almost impossible. It might greatly benefit the sender if the it has a mechanism to validate whether a super node has enough liquidity (balance) to route its payment without violating any privacy concerns in the network. Achieving this is hard because the local balance a node has with its immediate neighbors in a PCN is considered private information and the sender has to know whether the super node possess enough liquidity without comprising privacy.
@@
2) Channel verification in a PCN: To be a part of any PCN, nodes will open payment channels with other nodes in the network. Two nodes open a payment channel between them by posting a transaction to the blockchain. This transaction can be posted on the blockchain (as in UTXO based blockchains like BTC) or this transaction can be posted as a function call to an existing smart contract (as in blockchains that support smart contracts like ETH). In the most popular PCN, $\mathsf{LN}$, the channel verification procedure is very inefficient. It involves the node who intends to verify a channel to request the block in which the channel opening transaction has been included, verifying whether the transaction has been successfully executed by the validator/miner and finally verifying if the channel opening transaction corresponds to a 2-2 multi signature address on the blockchain. The verifier performing these steps is inefficient since all these steps will have already been performed by the miner. According to the current verification process, any node would need to perform all of these steps for each payment channel it has open in the network. Consider the following scenario: a node $j$ in the PCN has payment channels open with $m$ nodes in the PCN. Furthermore, assume that each neighbor opens $k$ new channels for each time epoch. For this example, assume the value of the time epoch is 1 block. This would mean that the total number of channels the node $j$ has to verify a total of $m*k$ channels for every block. This requires the node $j$ to access the blockchain a total of $m*k$ times to perform channel verification. Finding a way to do this without blockchain access and in a blockchain agnostic manner in a hard research challenge.
@@ VII-B Gaps in Tumblers
We will now describe the research gaps in tumblers. 1) Privacy against aborts: In a given a transaction, the tumbler should not be able to correctly identify the sender/receiver pair involved. Currently, there is only one construction [59] that guarantees this. However, this construction works only for blockchains that support Turing complete programming languages. Currently, there is no such construction for UTXO based blockchains that guarantees privacy against aborts. This is hard because if the sender aborts a transaction, the tumbler will know which receiver did not receive the payment. This is hard because, a tumbler mixes multiple transactions together to ensure unlikability among payments. Also, a tumbler should enforce atomicity so that the payment either goes through completely or does not go through at all. In the case of an abort, an adversary who controls the tumbler might be able to deduce which receiver amongst the ones involved did not receive a payment and will be able to link the corresponding transaction amount to that receiver.
@@
2) Time epochs: A tumbler should be able to facilitate payments without having a predefined time interval or epoch. Currently, the majority of payment channel hub constructions process payments in discrete epochs to increase the size of their anonymity sets. An anonymity set refers to a set of users (senders/receivers) whose identity is indistinguishable from each other. The size of this set is proportional to the privacy of sender/receiver in the transaction. When the interval between two consecutive time epochs is sufficiently large, it will promote more users to take part in the transactions within a time epoch, which improves privacy. Not having time epochs results in smaller anonymity sets and hence could possibly weaken privacy. The trade off is between privacy and flexibility.
@@
3) Dynamic corruption: A tumbler should ideally be resistant to dynamic corruption. This means the protocol should be resilient to parties deviating from their ideal execution during the beginning, progress, and ending of the protocol’s execution. Current tumbler constructions only provide support for static corruption, which may be unrealistic from a practical standpoint. Security against dynamic corruption is hard to achieve because the tumbler mixes several transactions together to preserve anonymity. Since the tumbler should not be able to link transaction amount to a sender/receiver pair (ideally), if the sender/receiver becomes malicious during the protocol’s execution, ideally, that transaction must be aborted to prevent financial loss to the honest party. However, the tumbler should not be able to link a transaction to a sender/receiver pair, making security against dynamic corruption a hard research problem.